# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ 5.2 Measured observables
Geometry correlation probe. A key requirement is that the geometry acts as a regulator rather than an additional critical field (“SR hypothesis”). To test this we measure a connected volume–volume correlator built from the cell areas $V(n)\equiv A(n)$. On each stored configuration we form $\delta V(n)=V(n)-\bar{V}$ with $\bar{V}=|\Lambda|^{-1}\sum_{m}V(m)$, and define
$C_{V}(m)=\frac{1}{|\Lambda|}\sum_{n\in\Lambda}\delta V(n)\,\delta V(n+m),$
where addition is modulo $L$ in each direction. We then radially bin this by graph distance $r(m)=\min(i,L-i)+\min(j,L-j)$ for $m=(i,j)$ and report
@@
Short–distance rotational probe. We measure an angle–binned finite–difference gradient observable: for a fixed list of short offsets $\Delta$ (nearest neighbours, diagonals, and a few second neighbours), we compute the minimal–image displacement $v$, its length $\ell=|v|$, and accumulate $(\Delta\phi/\ell)^{2}$ into $N_{\theta}=18$ bins in $0,\pi)$ by the orientation of $v$ (with $v\sim-v$). From the resulting profile $G(\theta)=\langle(\Delta\phi/\ell)^{2}\rangle(\theta)$ we form the anisotropy measures
$A_{\rm rms}=\frac{\sqrt{\langle(G(\theta)-\bar{G})^{2}\rangle_{\theta}}}{\bar{G}},\qquad A_{4}=\frac{\sqrt{C_{4}^{2}+S_{4}^{2}}}{\bar{G}},$
with $\bar{G}=\langle G(\theta)\rangle_{\theta}$, $C_{4}=\langle G(\theta)\cos 4\theta\rangle_{\theta}$, $S_{4}=\langle G(\theta)\sin 4\theta\rangle_{\theta}$.
@@ 5.3 Geometry sector diagnostics
Figure 3(a) shows a representative embedded mesh from the dynamical ensemble.
To quantify mesh fluctuations we show histograms of corner angles ($\theta$) in Fig. 3(b), edge lengths ($\ell$) in Fig. 3(c), cell areas ($A$) in Fig. 3(d), aspect ratios ($r$) in Fig. 3(e), and physical edge orientations in Fig. 3(f). In lattice units we find
@@
Geometry correlation lengths. A key requirement is that the geometry should act as a regulator rather than a critical field. We therefore measure two complementary correlators: (i) a displacement correlator based on u(n)=x(n)anu(n)=x(n)-an, with the configurationwise mean subtracted to remove the translation zero mode, and (ii) a connected volume–volume correlator $C_{V}(r)=\langle\delta V(0)\,\delta V(r)\rangle_{c}$ binned by graph distance. Both decay rapidly and are consistent with short–range geometry fluctuations, cf. Fig 3(g)3(h).
In plots we show $|C_{V}(r)|$ on a log scale and extract an effective decay length $\ell_{\rm geom}$ from an exponential fit over short to intermediate $r$. Across the volumes tested we find $\ell_{\rm geom}=O(1)$ in lattice units with no systematic growth with $L$ (e.g. $\ell_{\rm geom}\approx 1.6,\,0.74,\,0.86,\,0.65$ for L=16, 24, 32, 48L=16,\,24,\,32,\,48 at $\mu^{2}\simeq-1.34$), providing direct numerical support for the short-range geometry hypothesis (SR) in this two-dimensional setting.
@@ 5.4 Thermalisation and stability
Figures 4(a) and 4(b) show time-series diagnostics for representative baseline and dynamical runs. After the thermalisation window the actions and low moments fluctuate around stationary means with no visible long–term drift. In particular, Fig. 4(a) records the evolution of the minimum cell area and the maximum aspect ratio over the lattice; both remain stable throughout the production window, with no signs of slow drift or metastability, indicating that the geometry explores the admissible region rather than getting trapped near a boundary.
@@ 5.5 Rotational artefacts: short and intermediate scales
Angle–binned gradients. On the fixed square lattice $G(\theta)$ exhibits a clear fourfold pattern, while on the dynamical ensemble the modulation is strongly reduced (Fig. 4(e)). Table 1 summarises this via $A_{\rm rms}$ and $A_{4}$. We see that quantitatively, $A_{\rm rms}$ drops by $71\%$ (factor $\approx 3.5$) and $A_{4}$ drops by $55\%$ (factor $\approx 2.2$) relative to the baseline.
@@
Table 1: Anisotropy measures from the angle–binned gradient observable.
@@ 5.6 Universality and critical scaling
Finally, we check that coupling the scalar theory to the DLR geometry does not change its long–distance universality class. We simulate $L\in\{16,24,32,48\}$ and scan the bare mass parameter $\mu^{2}$ at fixed quartic coupling (here $\lambda=1$) using the protocol of Section 5.1. Figure 4(f) shows the Binder cumulant and $\xi/L$ as functions of $\mu^{2}$. Within our statistical precision, the baseline and dynamical regulators yield compatible crossing regions and scaling curves for both $U_{4}$ and $\xi/L$, supporting the conclusion that the coupled geometry does not change the long-distance universality class in this two-dimensional test. Both observables exhibit a common crossing region near $\mu_{c}^{2}\simeq-1.34$, consistent with a second–order transition and with the expected scale invariance at criticality. Deep in the ordered phase, finite–volume effects (including slow tunnelling between the two symmetry–broken sectors) can contaminate standard connected estimators; the crossing region itself remains sharp and well–defined.
We estimate integrated autocorrelation times for $M^{2}$ and observe the expected critical slowing down with $L$; at $\mu^{2}\approx-1.34$ we find $\tau_{M^{2}}\sim 1.7,\,3.1,\,5.3,\,7.6\times 10^{2}$ (for $L=16,24,32,48$), so production lengths and error analysis use binning/blocking with block size $\gg\tau_{M^{2}}$ and multiple independent seeds.
@@
Refer to caption (e) Angle-binned gradient profile $G(\theta)$.
@@ 6 Conclusions and outlook
In this paper we have introduced and analysed a dynamical–lattice regulator for Euclidean quantum field theories in which the abstract hypercubic topology is kept fixed, while the embedding of lattice sites into physical space is promoted to a dynamical geometry field $x:\Lambda\to\mathbb{R}^{d}$ integrated over in the path integral. The construction is local on the abstract lattice, preserves exact gauge invariance, and enjoys exact global $\mathrm{E}(d)$ invariance at finite lattice spacing. By treating $x$ as part of an extended local field multiplet $(x,U,\Phi)$ we have shown that Osterwalder–Schrader reflection positivity holds under mild assumptions on the geometry action $S_{x}[x]$, so that the usual Hamiltonian reconstruction goes through. Within a Symanzik effective–action framework, and conditional on the short-range geometry hypothesis (SR), geometry fluctuations generate only irrelevant operators in the continuum limit, and a one–loop calculation in scalar $\phi^{4}$ theory confirms that the $\beta$–function of the quartic coupling is unchanged. A simple Monte Carlo scheme on a two–dimensional torus suggests that the geometry sector can be simulated stably and that basic scalar observables agree with those on a static lattice at the same bare parameters.
From a structural point of view, the dynamical–lattice regulator defines a conservative middle ground between rigid hypercubic lattices and more radical fluctuating–topology approaches such as random lattices and dynamical triangulations. The abstract connectivity of the lattice is fixed, which makes reflection positivity, transfer–matrix constructions and reuse of existing lattice algorithms straightforward, while the embedding $x$ fluctuates with a local action inside the same path integral as the gauge and matter fields. In this sense the geometric randomness is intrinsic to the quantum Lagrangian rather than imposed by an external random mesh generator or interpreted as a discretised gravitational field. The framework is also reminiscent of Arbitrary Lagrangian–Eulerian formulations in numerical analysis, but with the crucial difference that the mesh motion is governed by a quantum action rather than by a prescribed deterministic algorithm.
@@ 6.1 Future directions
  • Systematic numerical studies. The exploratory scalar simulations outlined in Section 5 should be developed into a systematic study of rotational symmetry restoration, autocorrelations and finite–size effects, and extended to gauge theories in $d=3,4$. In particular, a quantitative comparison of anisotropy diagnostics and scaling violations between static and dynamical meshes would clarify the practical gains of the regulator.
@@
  • Geometry phase structure in $d=4$ and the short–range hypothesis. The viability of DLR as a regulator hinges on remaining in a massive/short–range geometry regime (SR), after fixing/quotienting the global $\mathrm{SE}(d)$ zero modes. While the $d=2$ scalar experiments provide evidence consistent with such a regime even for minimal dynamics, in $d=4$ this is not guaranteed: fluctuating-geometry ensembles may develop entropic instabilities or long-range geometric modes unless a suitable stiffness term is included. A priority for four–dimensional applications is therefore to map the geometry (or coupled gauge–geometry) phase diagram as a function of stiffness parameters and admissibility thresholds, using geometry correlators and autocorrelation times as diagnostics.
  • Gauge field diagnostics and roughening. Extending the code to include gauge fields would allow for the measurement of the static quark potential. A critical test of the regulator would be to verify the persistence of the Lüscher term (string vibration) in regimes where a static lattice would exhibit a roughening transition (freezing of the flux tube). This would serve as a robust physical confirmation that the dynamical mesh eliminates the grid anisotropy that pins topological objects.
  • Alternative geometry actions and admissible sets. Different functional forms of $S_{x}$ and different admissibility criteria (e.g. angle–based versus condition–number–based constraints) can be explored, both analytically and numerically, to understand how they influence mesh statistics and cutoff effects. It would be useful to identify simple classes of geometry actions that balance good geometric behaviour (no crumpling or clustering) with acceptable autocorrelation times.
@@
  • Fermions and chiral symmetry. In this first work we have not addressed the question of fermion doubling in detail. Since the dynamical mesh restores exact $\mathrm{E}(d)$ symmetry but does not change the local Dirac operator on the abstract lattice, one expects the usual no–go theorems to continue to apply. Nonetheless, it would be worthwhile to investigate systematically how standard fermion formulations (Wilson, staggered, overlap) behave in the presence of a fluctuating geometry. Specifically, one should quantify whether averaging over the mesh reduces rotational artefacts such as the taste-splitting in staggered fermions, effectively isotropizing the doubler spectrum even if their number is preserved.
@@
  • Other topologies and simplicial variants. While we have focused on a hypercubic abstract lattice for constructive and algorithmic reasons, a simplicial version of the dynamical–lattice regulator would be natural and would connect more directly to Regge calculus. One could also consider other topologies (e.g. manifolds with boundaries, or nontrivial spatial topology) as long as a suitable notion of admissible embeddings and OS reflection exists.
@@
We hope that the dynamical–lattice regulator developed here will serve as a useful addition to the toolbox of lattice field theory, both as a conceptual laboratory for studying the interplay between geometry and regularisation and, potentially, as a practical means of reducing rotational artefacts in nonperturbative simulations.

### v2

@@ 5.2 Measured observables
Geometry short-range diagnostics. A key requirement is that the geometry acts as a regulator rather than an additional critical field (“SR hypothesis”). To test this we measure a connected volume–volume correlator built from the cell areas $A(n)$. On each stored configuration we form $\delta A(n)=A(n)-\bar{A}$ with $\bar{A}=|\Lambda|^{-1}\sum_{m}A(m)$, and define the translation–averaged correlator
$C_{V}(m)\;=\;\frac{1}{|\Lambda|}\sum_{n\in\Lambda}\delta A(n)\,\delta A(n+m),$
where addition is modulo $L$ in each direction. The reported $C_{V}(m)$ is then the Monte–Carlo average of this quantity. We radially bin by the periodic graph ($\ell^{1}$) distance
$r(m)=\min(i,L-i)+\min(j,L-j)$
for $m=(i,j)$ and report
@@
Short–distance rotational probe. We measure an angle–binned finite–difference gradient observable: for a fixed list of short offsets $\Delta$ (nearest neighbours, diagonals, and a few second neighbours), we compute the minimal–image displacement $v$, its length $\ell=|v|$, and accumulate $(\Delta\phi/\ell)^{2}$ into $N_{\theta}=18$ bins in $0,\pi)$ by the orientation of $v$ (with $v\sim-v$). From the resulting profile $H(\theta)=\langle(\Delta\phi/\ell)^{2}\rangle(\theta)$ we form the anisotropy measures
$A_{\rm rms}=\frac{\sqrt{\langle(H(\theta)-\bar{H})^{2}\rangle_{\theta}}}{\bar{H}},\qquad A_{4}=\frac{\sqrt{C_{4}^{2}+S_{4}^{2}}}{\bar{H}},$
with $\bar{H}=\langle H(\theta)\rangle_{\theta}$, $C_{4}=\langle H(\theta)\cos 4\theta\rangle_{\theta}$, $S_{4}=\langle H(\theta)\sin 4\theta\rangle_{\theta}$.
@@ 5.3 Geometry sector diagnostics
Figure 3(a) shows a representative embedded mesh from the dynamical ensemble, illustrating that the geometry is genuinely fluctuating while remaining well inside the admissible region.
To quantify mesh fluctuations we show histograms of corner angles ($\theta$) in Fig. 3(b), edge lengths ($\ell$) in Fig. 3(c), cell areas ($A$) in Fig. 3(d), aspect ratios ($r$) in Fig. 3(e), and physical edge orientations in Fig. 3(f). In lattice units we find (mean $\pm$ standard deviation)
@@
Geometry correlation lengths. A key requirement is that the geometry should act as a regulator rather than a critical field. We therefore measure two complementary correlators: (i) a displacement correlator based on u(n)=x(n)anu(n)=x(n)-an, with the configurationwise mean subtracted to remove the translation zero mode, and (ii) the connected volume–volume correlator $C_{V}(r)$ defined in §5.2, binned by graph distance. Both decay rapidly and are consistent with short–range geometry fluctuations, cf. Fig 3(g)3(h).
In plots we show $|C_{V}(r)|$ on a log scale and extract an effective decay length $\ell_{\rm geom}$ from an exponential fit over short to intermediate $r$. Across the volumes tested, the fitted decay length stays $O(1)$ in lattice units with no systematic growth in $L$ (e.g. $\ell_{\rm geom}\approx 1.34,\,0.99,\,0.54,\,0.92$ for L=16, 24, 32, 48L=16,\,24,\,32,\,48 at $\mu^{2}\simeq-1.40$). This provides direct numerical support for (SR) in this two–dimensional test.
@@ 5.4 Thermalisation and stability
Figures 4(a) and 4(b) show time-series diagnostics for representative baseline and dynamical runs. After the thermalisation window the actions and low moments fluctuate around stationary means with no visible long–term drift. Figure 4(a) records the evolution of the minimum cell area and the maximum aspect ratio over the lattice; both remain stable throughout the production window, with no signs of slow drift or metastability, indicating that the geometry explores the admissible region rather than getting trapped near a boundary.
@@ 5.5 Rotational artefacts: short and intermediate scales
Angle–binned gradients. On the fixed square lattice $H(\theta)$ exhibits a clear fourfold pattern, while on the dynamical ensemble the modulation is strongly reduced (Fig. 4(e)). Table 2 summarises this via $A_{\rm rms}$ and $A_{4}$. Quantitatively, $A_{\rm rms}$ drops from $0.42$ to $0.12$ (a factor $\approx 3.5$), while $A_{4}$ drops from $0.17$ to $0.08$ (a factor $\approx 2.2$), indicating substantial suppression of the dominant fourfold harmonic in this short–distance diagnostic.
@@
Table 2: Anisotropy measures from the angle–binned gradient observable ($\mu^{2}=-1.34$, $L=48$).
@@ 5.6 Universality and critical scaling
Finally, we check that coupling the scalar theory to the DLR geometry does not change its long–distance universality class. We simulate $L\in\{16,24,32,48\}$ and scan the bare mass parameter $\mu^{2}$ at fixed quartic coupling (here $\lambda=1$) using the protocol of Section 5.1. Figure 4(f) shows the Binder cumulant and $\xi/L$ as functions of $\mu^{2}$ for the dynamical ensemble. Both observables exhibit a common crossing region near $\mu_{c}^{2}\simeq-1.40$, as expected for a second–order transition. The Binder value at the crossing is consistent with the 2D Ising universality class; since the baseline square–lattice $\phi^{4}$ theory is known to lie in the same class, this supports the conclusion that the coupled geometry does not modify the long–distance universality in this two–dimensional test.
We estimate integrated autocorrelation times for $M^{2}$ and observe the expected critical slowing down with $L$; at $\mu^{2}\approx-1.40$ we find $\tau_{M^{2}}\approx(1.4,\,4.0,\,5.2,\,8.5)\times 10^{2}$ sweeps for $L=16,24,32,48$. Accordingly, error bars are obtained with blocking/binning using block sizes $\gg\tau_{M^{2}}$ and by averaging over multiple independent seeds. As a basic check on the geometry sector, we also monitored autocorrelations of simple global geometry summaries (e.g. $\sum_{n}A(n)$ and $\sum_{n}S_{x}(n)$) and found no anomalously slow mode in the runs shown.
@@
Refer to caption (e) Angle-binned gradient profile $H(\theta)$.
@@ 6 Conclusions and outlook
In this paper we established that the dynamical–lattice regulator (DLR) provides a local, gauge-invariant cutoff that is exactly $\mathrm{SE}(d)$–covariant at finite lattice spacing, and admits Osterwalder–Schrader reflection positivity (hence a standard transfer-matrix reconstruction) for a broad class of coupled geometry–field actions. Assuming the short-range geometry hypothesis (SR), integrating out $x$ at fixed $a$ yields a local Symanzik effective action in which geometry fluctuations generate only $\mathrm{SO}(d)$–invariant irrelevant operators (and finite renormalisations), a mechanism checked explicitly at one loop in scalar $\phi^{4}$. Our $d=2$ Monte Carlo tests provide proof-of-concept evidence for stable short-range geometry near matter criticality, reduced rotational artefacts at finite cutoff, and matching critical scaling with the baseline theory.
From the lattice–QCD viewpoint, DLR may be regarded as an annealed geometric averaging in coordinate space that preserves the transfer–matrix/reflection–positivity framework, complementary to field–space smoothing techniques (smearing, gradient flow) that improve operators/observables but do not on their own define a new reflection–positive cutoff. Structurally, DLR also occupies a conservative middle ground between rigid hypercubic regulators and fluctuating–topology approaches such as random lattices and dynamical triangulations: the abstract connectivity is fixed (so reflection positivity and standard lattice algorithms remain accessible), while the embedding fluctuates with a local action inside the same path integral as the gauge and matter fields.
@@ 6.1 Future directions
  • Systematic numerical studies. The exploratory scalar simulations in Section 5 should be developed into a systematic study of rotational symmetry restoration, autocorrelations and finite–size effects, and extended to gauge theories in $d=3,4$. In the gauge sector this should include standard benchmarks such as plaquette distributions, gradient–flow scales, the static quark potential and string tension (including the Lüscher term), as well as diagnostics of topological objects and topological–charge autocorrelations (topology freezing) and their sensitivity to grid artefacts. A quantitative comparison of anisotropy diagnostics and scaling violations between static and dynamical meshes would clarify the practical gains of the regulator.
@@
  • Geometry phase structure in $4d$ and the short–range hypothesis. The viability of DLR as a regulator hinges on an operating window in which the geometry sector remains short-range (SR) after fixing the global $\mathrm{SE}(d)$ zero modes: connectivity is fixed and nondegenerate, yet local orientations remain disordered. While the $d=2$ scalar experiments provide evidence consistent with such a regime even for a minimal geometry action, in $4d$ this is not guaranteed: competing entropic regimes (e.g. crystallisation/orientational ordering or, at the opposite extreme, crumpling–type instabilities familiar from random geometry models) may appear unless suitable stiffness terms and admissibility thresholds are chosen. A priority for $4d$ applications is therefore to map the geometry phase diagram as a function of these parameters, using geometry correlators and autocorrelation times as diagnostics.
@@
  • Fermions and chiral symmetry. In this first work we have not addressed the question of fermion doubling in detail. Since the dynamical mesh restores exact $\mathrm{SE}(d)$ symmetry but does not change the local Dirac operator on the abstract lattice, one expects the usual no–go theorems to continue to apply. Nonetheless, it would be worthwhile to investigate systematically how standard fermion formulations (Wilson, staggered, overlap) behave in the presence of a fluctuating geometry, and to quantify whether averaging over the mesh reduces rotational artefacts such as taste splitting in staggered fermions, effectively isotropising the doubler spectrum even if their number is preserved.
  • Quenched geometry and “quantum adaptivity”. Besides the annealed formulation studied here, it would be interesting to consider a quenched variant in which the abstract hypercubic connectivity is fixed and the embedding $x(n)=a\,n+\eta(n)$ is sampled independently from an $\mathrm{SE}(d)$–invariant short–range law, after which gauge and matter fields are simulated conditionally on $x$ and observables are averaged over the geometry ensemble. This quenched perspective removes geometric backreaction and may therefore be more amenable to analytic control. Conversely, in the annealed theory the fields can bias the geometry measure; it is an interesting possibility that this backreaction acts as a kind of quantum adaptivity, with the auxiliary embedding degrees of freedom tending to adjust locally in response to field configurations, potentially reducing certain cutoff artefacts.
  • Hamiltonian formulation. An important direction is to develop a Kogut–Susskind–type Hamiltonian version of DLR. Given Osterwalder–Schrader reflection positivity and the transfer matrix interpretation of the Euclidean theory, one expects a canonical description in which geometry variables appear as additional degrees of freedom on time slices and modify the electric and magnetic terms only via local geometry–dependent couplings.
@@
  • Other topologies and simplicial variants. While we have focused on a hypercubic abstract lattice for constructive and algorithmic reasons, a simplicial version of DLR would be natural and would connect more directly to Regge calculus. One could also consider other topologies (e.g. manifolds with boundaries, or nontrivial spatial topology) as long as a suitable notion of admissible embeddings and OS reflection exists.
@@
We view the present work as a starting point, and we invite further numerical and analytical tests of DLR, in particular in $d=4$ where establishing (SR) and benchmarking rotational artefacts against standard improvement and flow–based techniques are decisive.